OpenStudy (anonymous):

For the given statement Pn, write the statements P1, Pk, and Pk+1.

OpenStudy (anonymous):

whats up with the 3.2 ?

OpenStudy (anonymous):

woops sorry. just 2. no 3.

OpenStudy (anonymous):

2 + 4 + 8 + . . . + 2n = 2^(n+1) - 2

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@amistre64 @badi @Callisto @experimentX @Hero @phi @.Sam.

OpenStudy (anonymous):

P(1) replace \(n\) by \(1\)

OpenStudy (anonymous):

can you show me how to work this out.

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

if you put \(n=1\) you get \(2=2^{1+1}-2\)

OpenStudy (anonymous):

where i saw an \(n\) i replaced it by \(1\) this is clearly true, since \[2=2^2-2=4-2=2\] is true

OpenStudy (anonymous):

\(P_k\) is also a replacement, replace \(n\) by \(k\) i.e write exactly what you see with \(k\) instead of \(n\)

OpenStudy (anonymous):

\[P(n):2 + 4 + 8 + . . . + 2n = 2^{n+1} - 2\] \[P(k):2 + 4 + 8 + . . . + 2k = 2^{k+1} - 2\]

OpenStudy (anonymous):

and for \(P(k+1)\) replace \(n\) by \(k+1\)

OpenStudy (anonymous):

ty :)